# Invisible Servoing: a Visual Servoing Approach with Return-Conditioned Latent Diffusion

### v1

@@ 5.2 Receding horizon control
5.3 Real-world experiment
Figure 7: Closed-loop trajectory of the hexarotor in the flight lab simulation environment.
We have mounted an Intel Realsense D435i on a multi-rotor UAV. The onboard Raspberry PI 5 is controlling the multi-rotor [^4] https://git.openrobots.org/projects/telekyb3 and runs the Realsense driver [^5] https://git.openrobots.org/projects/realsense-genom3 that publishes via ROS the raw image with the same resolution $340\times 180$ as the one in the dataset. A ground station PC runs the node that subscribes to the camera topic, samples the DDPM with inpainting conditioning on $R_{est}$, and publishes the set of waypoints obtained integrating the velocity control actions. The waypoints are then tracked by the motion controller onboard. An end-to-end approach to directly publish low-level controls is left as future work.
The real-world tests show that the diffusion model trained in simulation can successfully inpaint a trajectory between the initial and the final desired image in the real world. The attached video and Fig. 1 show that the proposed receding horizon approach allowed reaching the visual target even if it is initially out of the view. It is worth highlighting that we do not rely on any real-world data to train the CM-VAE and the DDPM, but only simulated data. Learning a regularized latent space and planning in it are the essential ingredients for transferring the learned planner to the real-world.
@@ 6 Conclusion
In this paper, we presented a novel VS approach based on latent DDPM, capable of handling target-invisible views. It relies on coupling a CM-VAE for the latent representation and DDPMs for planning latent space trajectories. Additionally, we introduced a heuristic to estimate the desired return to be conditioned on that drastically improves the generated trajectories by the DDPM. This work used only simulation data for training, yet achieved successful real-world experiments, demonstrating the effectiveness of our approach in seamless sim-to-real transfer.
Future work includes speeding up the sampling process, to allow for a full end-to-end approach. The following strategies will be considered: (i) using more powerful hardware, e.g., Nvidia Jetson, and reducing the diffusion steps using (ii) warm-start strategies, i.e., exploiting previously-generated trajectories as initial guesses for the DDPM to denoise the next trajectory, and (iii) leveraging the research on novel and computationally efficient DDPMs, such as the Denoising Diffusion Implicit Models [29].
@@ Acknowledgments
The authors thank Quentin Sablé (University of Twente) for the technical support, Anthony Mallet (LAAS-CNRS) for the camera driver software support, along with all the contributors to the Aerial Robotics Testbed project [^6] https://git.openrobots.org/projects/art.

### v2

@@ 6 Conclusion
In this paper, we presented a novel VS approach based on latent DDPM, capable of handling target-invisible views. It relies on coupling a CM-VAE for the latent representation and DDPMs for planning latent space trajectories. Additionally, we introduced a heuristic to estimate the desired return to be conditioned on that drastically improves the generated trajectories by the DDPM.
Future work includes speeding up the sampling process, to allow for a full end-to-end approach. The following strategies will be considered: (i) using more powerful hardware, e.g., Nvidia Jetson, and reducing the diffusion steps using (ii) warm-start strategies, i.e., exploiting previously-generated trajectories as initial guesses for the DDPM to denoise the next trajectory, and (iii) leveraging the research on novel and computationally efficient DDPMs, such as the Denoising Diffusion Implicit Models [32]. Also, the approach will be experimentally validated in real-world and more complex demonstration scenarios.
@@ Acknowledgments
The authors thank Anthony Mallet (LAAS-CNRS) for the camera driver software support, along with all the contributors to the Aerial Robotics Testbed project [^4] https://git.openrobots.org/projects/art.